OpenStudy (anonymous):

Need Help!!! A spotlight on the ground shines on a wall 12 m away. If a man 2 m tall walks from the spotlight toward the building at a speed of 2.3 m/s, how fast is the length of his shadow on the building decreasing when he is 4 m from the building? (Round your answer to one decimal place.)

OpenStudy (anonymous):

I got -0.8625 and rounded to -0.9. Don't know why it's incorrect

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{2}{x}=\frac{s}{12}\] and is what i get as a start

OpenStudy (anonymous):

making \[s=\frac{24}{x}\] you want \(s'\) taking derivative we get \[s'=-\frac{24}{x^2}x'\] you know \(x'=2.3, x=8\) substitution should work

OpenStudy (anonymous):

how to you like that, i got the same answer as you

OpenStudy (anonymous):

i don't know why it is incorrect either, whatever it is we made the same mistake or, more likely, whatever system you are using wants more or fewer decimals. who knows?